Find the equilibrium of income level and interest rate and the level of C, I, Mt and M2 in equilibrium value: C=81²+0.6Y, I= 120-150i, Ma= 275,Mt= 0.1Y and M2= 240-250i
1
Expert's answer
2019-04-04T09:20:31-0400

If

"Ms = Md"

and

"Y = C + I,"

then:

0.1Y + 240 - 250i = 275 -> Y = 35 + 250i,

81 + 0.6Y + 120 - 150i = Y - > Y = 502.5 - 375i,

35 + 250i = 502.5 - 375i,

625i = 467.5,

i = 0.75,

Y = 35 + 250*0.75 = 222.5.